kaylak:

calc help like maybe 15 questions

Falconmaster:

@Shadow @dude halp dis grill

jhonyy9:

??? where are ?

Kamauri:

i dont see any

Shadow:

@kaylak please post the questions when you can.

kaylak:

kaylak:

@Vocaloid @Shadow

kaylak:

@Vocaloid

Vocaloid:

for 1) inflection points you have to take the second derivative and see if the sign changes from + to - anywhere (graphing is a shortcut method to accomplish this)

Vocaloid:

3) another concavity question, take the second derivative and identify where its positive

Vocaloid:

4) if 0 is the lowest output value this function can go, then it's true; otherwise false

kaylak:

is 1 true?? I will have to look again

Vocaloid:

two pieces of evidence point to false: 1. if you look at the graph of ln(5x) it's concave down everywhere (so it never changes concavity and thus has no inflection point) 2. the second derivative is -1/x^2 which is always negative

kaylak:

ah okay sorry Im not on a lot to work it out busy cleaning let me see if I can determine next one and tonight is just 8 questions

kaylak:

Is 2 true?

Vocaloid:

yeah

kaylak:

3 is a ?

Vocaloid:

yeah that's what i got too

kaylak:

4 is it true

Vocaloid:

yes

kaylak:

kaylak:

last 4 for tonight

kaylak:

15 for 1? let me check

kaylak:

@Vocaloid

kaylak:

it's -7 at least that's what I got

kaylak:

true?

Vocaloid:

I don't remember how to do 1) using calculus but just graphing it gives the maximum value at x = 5 which gives you 11 when you plug it into the function

kaylak:

okay

Vocaloid:

same logic w/ 2) I ended up graphing it and found that x = -pi/2 is not a relative minimum

kaylak:

okay so you are confident on 5 of 6?

Vocaloid:

for 3) I drew the cardboard piece out: 

Vocaloid:

to save some time you can either use a calculator to find the maximum or take the derivative either way the maximum volume ends up being 24 which would fall between the 20-30 range

kaylak:

okay cool

Vocaloid:

for the last one, you just have to multiply each pair of numbers and see which one gives the smallest product so 39*1 = 39 15*25 = 375 10*30 = 300 20*20 = 400 making 39 and 1 the best choice

Vocaloid:

as a general rule, if the two factors have a set sum, the farther the factors are, the smaller the product will be; the closer the two factors are, the higher the product will be

kaylak:

I didn't even think of that now I feel slow ty so much

kaylak:

are you confident with these answers? so I can study this

Vocaloid:

I feel pretty good but a 2nd opinion wouldn't hurt

kaylak:

100% ty!